OpenStudy (anonymous):

ABCD is a rectangle. If AC =3y/5 and BD = 3(y – 4), find the length of each diagonal. a.−3, −3 b.5, 5 c. 3,3 d 16,16

jimthompson5910 (jim_thompson5910):

In any rectangle, the diagonals are congruent

jimthompson5910 (jim_thompson5910):

so AC = BD 3y/5 = 3(y-4) 3y = 5*3(y-4) 3y = 15(y-4) 3y = 15y - 60 Keep going to solve for y

jimthompson5910 (jim_thompson5910):

once you have the value of y, use it to find AC or BD